OpenStudy (chazz16):

Help. MEDAL WILL BE AWARDED! Find The Missing Factor. Write Your Answer In Exponential Form. 9^2 = 9^4 * ____?_____

OpenStudy (trojanpoem):

9^2 = 9^4 * 9^x x+ 4 =2 x = -2

OpenStudy (trojanpoem):

You could just eye-spot it , but the steps for more complicated examples.